# Occlusion-Aware Ground Target Tracking by a Dubins Vehicle Using Visibility Volumes

### v1

@@ 6.1 Vector Field Guidance for the Desired Heading
where the positive solution is counter-clockwise rotation and the negative solution is clockwise rotation. Expanding (60) with (65), (72), (73), and (74) results in
$\bm{u}$ $=\left(-\Phi^{\prime}+\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\right)\bm{e}_{r}$
±$\pm\left(\sqrt{v^{2}-\left(-\Phi^{\prime}+\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\right)^{2}}\right)\bm{e}_{\theta}\;.$ (75)
The vector field (75) consists of three parts visualized in Fig. 6: (i) $-\Phi^{\prime}\bm{e}_{r}$ radial attraction, (ii) $(\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r})\bm{e}_{r}$ the feed-forward component, and (iii) circulation about the POI $\left(\sqrt{v^{2}-\left(-\Phi^{\prime}+\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\right)^{2}}\right)\bm{e}_{\theta}$.
Figure 6: The three parts of a time-varying circulation vector field in (60) for counter-clockwise circulation. Left: The attraction component of the vector field $\Phi\nabla V$ pushes initial conditions onto the orbit. Center: The component of the vector field $\bm{M}^{-1}\dot{\bm{\sigma}}$ that forces the UAV to follow a POI that is moving diagonally across the page from the top right towards the bottom left. Right: The circulation component of the vector field $H\wedge_{i=0}^{n-1}\nabla a_{i}$ forces the initial conditions to circulate about the POI.
The combination of the radial attraction and the feed-forward terms $(-\Phi^{\prime}+\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r})\bm{e}_{r}$ are shown in Fig. 7 to compare the magnitude of the two radial fields.
Figure 7: The combination of the radial attract term and the feed-forward term.
Trajectories of a particle moving in $\mathbb{R}^{2}$ using the controller $\bm{u}$ are visualized in Fig. 8. It can be seen that the velocity streamlines do not arrive tangentially at the orbit because the orbit is translating to the bottom right corner—the UAV will not remain on the orbit in the next time step if the streamlines are tangent to the orbit.
Figure 8: The vector field $\bm{u}$ for a circular orbit (black) of $R=10$ m with POI (black) moving with speed $v=5$ m/s with direction $\gamma=\frac{-3\pi}{4}$, the value $\beta=0.1$, and the UAV’s speed is $v=10$ m/s. The constant magnitude vector field for counter-clockwise circulation is shown in gray. The left plot shows time $t=0$ and the right plot shows time $t=1$ s.
@@
$\begin{split}\bm{u}&=\left(\left(-v+v_{g}+\left|\dot{R}\right|\right)\frac{2}{\pi}\arctan(\beta(r-R))+\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R}\right)\bm{e}_{r}\\ &\pm\left(\sqrt{v^{2}-\left(\left(-v+v_{g}+\left|\dot{R}\right|\right)\frac{2}{\pi}\arctan\left(\beta(r-R)\right)+\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R}\right)^{2}}\right)\bm{e}_{\theta}\;.\end{split}$ (76)
This controller builds upon the framework in [54] by finding a specific implementation of the generalized controller for a constant velocity single integrator model. To find the implementation of the controller, this work found a function $a_{1}$, a function $\Phi$, and a function $H$ that created a constant magnitude vector field that is tailored towards tracking the desired reference orbits introduced in Sec. 5.
@@ 6.2 Heading Rate
The vector field (76) can also be written in terms of polar coordinates $r$ and $\theta$ and related to the rectangular coordinates with the rotation,
$\begin{bmatrix}u_{x}\\ u_{y}\end{bmatrix}$ $=u_{r}\bm{e}_{r}+u_{\theta}\bm{e}_{\theta}$
$=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}u_{r}\\ u_{\theta}\end{bmatrix}\;.$ (77)
Since the magnitude of the vector field is constant and the same as the Dubins vehicle speed, (32) can be rewritten in rectangular coordinates $x$ and $y$ as
@@
which is a function of the current $xy$ position and time $t$. If the the UAV starts with initial condition $\bm{q}_{0}=[q_{x},q_{y},\psi_{d}([q_{x},q_{y}]^{\rm T},0)]^{\rm T}$ then a vehicle with an unbounded turn rate can flow along the vector field $\bm{u}$ with the dynamics
$\begin{bmatrix}\dot{q}_{x}\\ \dot{q}_{y}\\ \dot{q}_{\psi}\end{bmatrix}=\begin{bmatrix}v\cos q_{\psi}\\ v\sin q_{\psi}\\ \dot{\psi}_{d}\end{bmatrix}\;,$ (80)
where the rate of change of the direction $\psi_{d}$ is
$\dot{\psi}_{d}=\frac{\dot{u}_{x}u_{y}-\dot{u}_{y}u_{x}}{u_{x}^{2}+u_{y}^{2}}=v\kappa_{s}\;,$
the expression of curvature from (4). Lemma 2 derives an expression for the curvature of a constant velocity curve in polar coordinates. By using the substitution $\dot{\bm{x}}_{\mathcal{I}}=\dot{\bm{\xi}}=\bm{u}$ with equation (21) from Lemma 2,
@@
where $\ddot{\bm{g}}=0$ since the POI moves in a straight line at a constant speed, $\ddot{R}=0$ since the radius changes linearly. The time derivative of (72) is
@@
$=\frac{(\dot{\xi}_{x}-\dot{g}_{x})(\xi_{x}-g_{x})+(\dot{\xi}_{y}-\dot{g}_{y})(\xi_{y}-g_{y})}{\sqrt{(\xi_{x}-g_{x})^{2}+(\xi_{y}-g_{y})^{2}}}\;,$ (84)
which can be simplified by substituting in (59), (12), and the polar direction $\bm{e}_{r}$ to give
$\dot{r}$ $=(\bm{u}-\dot{\bm{g}})\cdot\bm{e}_{r}$
$=u_{r}-\dot{\bm{g}}\cdot\bm{e}_{r}\;.$
Substituting (75) further simplifies this expression to
@@
$\dot{\Phi}^{\prime}$ $=\left(v-v_{g}-\left|\dot{R}\right|\right)\frac{2}{\pi}\frac{\beta((-\Phi^{\prime}+\dot{R})-\dot{R})}{1+\beta^{2}(r-R)^{2}}$
$=\left(v-v_{g}-\left|\dot{R}\right|\right)\frac{2}{\pi}\frac{\beta(-[\left(v-v_{g}-\left|\dot{R}\right|\right)\frac{2}{\pi}\arctan\left(\beta(r-R)\right)])}{1+\beta^{2}(r-R)^{2}}$
@@
$=\frac{\frac{(\dot{\xi}_{y}-\dot{g}_{y})(\xi_{x}-g_{x})-(\dot{\xi}_{x}-\dot{g}_{x})(\xi_{y}-g_{y})}{(\xi_{x}-g_{x})^{2}}}{1+\left(\frac{\xi_{y}-g_{y}}{\xi_{x}-g_{x}}\right)^{2}}$
$=\frac{(\dot{\xi}_{y}-\dot{g}_{y})(\xi_{x}-g_{x})-(\dot{\xi}_{x}-\dot{g}_{x})(\xi_{y}-g_{y})}{(\xi_{x}-g_{y})^{2}+(\xi_{y}-g_{y})^{2}}$
@@
$\dot{\theta}$ $=\frac{(\bm{u}-\dot{\bm{g}})\cdot\bm{e}_{\theta}}{r}$
$=\frac{u_{\theta}-\dot{\bm{g}}\cdot\bm{e}_{\theta}}{r}\;.$ (87)
@@
$\dot{\psi}_{d}$ $=\frac{\dot{\Phi}^{\prime}-\dot{\theta}\dot{\bm{g}}\cdot{{\bm{e}}_{\theta}}}{u_{\theta}}+\dot{\theta}$
$=\frac{\dot{\Phi}^{\prime}}{u_{\theta}}+\dot{\theta}\left(\frac{u_{\theta}-\dot{\bm{g}}\cdot\bm{e}_{\theta}}{u_{\theta}}\right)$
$=\frac{\dot{\Phi}^{\prime}}{u_{\theta}}+\dot{\theta}\frac{r}{u_{\theta}}\left(\frac{u_{\theta}-\dot{\bm{g}}\cdot\bm{e}_{\theta}}{r}\right)\;.$
@@
The value of $\dot{\psi}_{d}$ at a range of inputs is shown in Fig. 9 where the UAV’s speed is $10$ m/s, the POI moves with speed $5$ m/s in direction $\frac{-3\pi}{4}$, the gain $\beta=0.1$, and $t=0$ s. The value of $\dot{\psi}_{d}$ is small for all values outside of the orbit’s radius $10$ m, but there is a singularity where $\dot{\psi}_{d}$ goes to infinity at $r=0$ m. For large values of $\beta$ the sign is not strictly positive or negative.
Figure 9: A surface plot and contour plots of $\dot{\psi}_{d}$ where UAV’s speed is $v=10$ m/s, the POI moves with speed $v_{g}=5$ m/s in direction $\frac{-3\pi}{4}$, the gain $\beta=0.1$, $t=0$ s, and the circulation direction is counter-clockwise. Both plots show the orbit with $R=10$ m as a dashed line. The middle of the plot is cropped to $\dot{\psi}_{d}=4$ but extends infinitely.
@@ 6.3 Attraction to the Orbit
If the UAV is initialized with a heading $q_{\psi,0}\in\mathbb{S}$ that is not aligned with the vector field ${\bm{u}}$ then a controller is needed to drive it towards $\psi_{d}$.
Remark 6.
The system (1) can reach a configuration $\bm{q}_{1}=[q_{1,x},q_{1,y},\psi_{d}([q_{1,x},q_{1,y}]^{\rm T},t_{1})]^{\rm T}$ on the vector field (76) using a Dubins path at a time $t\geq t_{\rm max}$ where
$t_{\rm max}=\frac{1}{v}\left(d+\zeta\pi r_{\rm min}\right)\;.$
[…]
The above remark helps illustrate that the problem of converging to the vector field is well posed. In principle, a Dubins based path planning scheme could be developed to construct this initial transition path to the vector field. However, such an approach may be computationally intensive. In the following we propose a feedback-based strategy instead. The following steering controller is proposed:
$u_{\psi}$ $=k_{\psi}\left(\psi_{d}-q_{\psi}\right)+\dot{\psi}_{d}$
$+\left(\left(\bm{g}\cdot\bm{e}_{r}+\dot{R}\right)\frac{1-\cos(q_{\psi}-\psi_{d})}{q_{\psi}-\psi_{d}}\right.$
$\left.+\frac{u_{\theta}\sin(q_{\psi}-\psi_{d})}{q_{\psi}-\psi_{d}}\right)\frac{\beta\frac{2}{\pi}\arctan\left(\beta(r-R)\right)}{1+\beta^{2}(r-R)^{2}}\;,$ (89)
where $\dot{\psi}_{d}$ is the rate of change of the heading for a trajectory flowing along the vector field $\bm{u}$, $k_{\psi}$ is a positive proportional gain, and the final term (derived in the following lemma) is required for Lyapunov-like stability [65].
@@
From [65, p. 125], a corollary of Barbalat’s lemma is the following. If a scalar function $V(\bm{q},t)$ satisfies, (i) $V(\bm{q},t)$ is lower bounded, (ii) $\dot{V}(\bm{q},t)$ is negative semi-definite, and (iii) $\dot{V}(\bm{q},t)$ is uniformly continuous in time, then $\dot{V}(\bm{q},t)\rightarrow 0$ as $t\rightarrow\infty$.
@@
where $\tilde{r}=r-R$ is the error between the radial position of the UAV with respect to the moving point $\bm{g}$ and $\tilde{\psi}=q_{\psi}-\psi_{d}$ is the error between the current heading of the UAV and the desired UAV heading where $\bm{\xi}=[q_{x},q_{y}]^{\rm T}$. The angular error $\tilde{\psi}\in\mathbb{S}$ is the smaller of two the distances from $\psi_{d}$ to $q_{\psi}$, traveling clockwise or counter-clockwise in $\mathbb{S}$. The function (91) has a lower bound of zero when $\tilde{\psi}=0$ and $\tilde{r}=0$. The total derivative of (91) is
@@
$\dot{\tilde{r}}$ $=\dot{r}-\dot{R}$
$=\frac{(\dot{q}_{x}-\dot{g}_{x})(q_{x}-g_{x})+(\dot{q}_{y}-\dot{g}_{y})(q_{y}-g_{y})}{\sqrt{(q_{x}-x)^{2}+(q_{y}-g_{y})^{2}}}-\dot{R}$
$=\left(\begin{bmatrix}\dot{q}_{x}\\ \dot{q}_{y}\end{bmatrix}-\begin{bmatrix}\dot{g}_{x}\\ \dot{g}_{y}\end{bmatrix}\right)\cdot\bm{e}_{r}-\dot{R}$
$=v\cos q_{\psi}\cos\theta+v\sin q_{\psi}\sin\theta-\dot{\bm{g}}\cdot\bm{e}_{r}-\dot{R}\;.$ (93)
@@
$\dot{\tilde{r}}$ $=v\cos(\psi_{d}+\tilde{\psi})\cos\theta+v\sin(\psi_{d}+\tilde{\psi})\sin\theta-\dot{\bm{g}}\cdot\bm{e}_{r}-\dot{R}$
$=v\cos\theta(\cos\psi_{d}\cos\tilde{\psi}-\sin\psi_{d}\sin\tilde{\psi})$
$+v\sin\theta(\sin\psi_{d}\cos\tilde{\psi}+\sin\tilde{\psi}\cos\psi_{d})-\dot{\bm{g}}\cdot\bm{e}_{r}-\dot{R}\;.$ (94)
@@
$\dot{\tilde{r}}$ $=u_{x}\cos{\theta}\cos\tilde{\psi}-u_{y}\cos\theta\sin\tilde{\psi}$
$+u_{y}\sin\theta\cos\tilde{\psi}+u_{x}\sin\theta\sin\tilde{\psi}-\dot{\bm{g}}\cdot\bm{e}_{r}-\dot{R}$
$=\cos\tilde{\psi}(u_{x}\cos\theta+u_{y}\sin\theta)$
$-\sin\tilde{\psi}(-u_{x}\sin\theta+u_{y}\cos\theta)-\dot{\bm{g}}\cdot\bm{e}_{r}-\dot{R}\;.$ (95)
@@
$\dot{V}_{2}$ $=-k_{\psi}\tilde{\psi}^{2}-\left[\left(v-v_{g}-\left|\dot{R}\right|\right)\beta\frac{2}{\pi}\arctan{\beta\tilde{r}}\right.$
$+\bm{g}\cdot\bm{e}_{r}+\dot{R}-\bm{g}\cdot\bm{e}_{r}-\dot{R}\bigg{]}\cos\tilde{\psi}\frac{\beta\frac{2}{\pi}\arctan{\beta\tilde{r}}}{1+\beta^{2}\tilde{r}^{2}}$
$=-k_{\psi}\tilde{\psi}^{2}-\frac{\left(v-v_{g}-\left|\dot{R}\right|\right)\cos(\tilde{\psi})\beta\frac{4}{\pi^{2}}\arctan^{2}{\beta\tilde{r}}}{1+\beta^{2}\tilde{r}^{2}}\;.$ (100)
@@
denote the coefficient appearing in (100). The bounds of $\epsilon$ were found using a numerical method to be $0\leq\epsilon\leq\epsilon_{\rm max}$ where $\epsilon_{\rm max}$ is the solution to the equation $01-x\arctan{x}=$ 0, which is approximately $\epsilon_{\rm max}\approx 0.31483$. The function $\epsilon$ can be rewritten as
@@
$=\lim_{x\rightarrow\pm\infty}\frac{x}{2x(1+x^{2})}+\frac{\arctan{x}}{2x}$
$=\lim_{x\rightarrow\pm\infty}\frac{1}{2(1+x^{2})+4x^{2}}+\frac{1}{2(1+x^{2})}$
$=0\;.$
@@
$k_{\psi}>v\beta\frac{4}{\pi^{4}}\epsilon_{\rm max}\geq\frac{\left(v-v_{g}-\left|\dot{R}\right|\right)\beta\frac{4}{\pi^{4}}\arctan^{2}{\beta\tilde{r}}}{1+\beta^{2}\tilde{r}^{2}}$ (105)
Similarly, substituting $\tilde{\psi}=\pm\frac{\pi}{2}$ into (100), the inequality $\dot{V}<0$ leads to
$k_{\psi}>0\;.$ (106)
To find the local extrema, find where the partial derivative is zero:
$\frac{\partial}{\partial\tilde{\psi}}\dot{V}_{2}$ $=0$
00 0 $=-2k_{\psi}\tilde{\psi}+\frac{\left(v-v_{g}-\left|\dot{R}\right|\right)\sin{\tilde{\psi}}\beta\frac{4}{\pi^{2}}\arctan^{2}{\beta\tilde{r}}}{1+\beta^{2}\tilde{r}^{2}}$
@@
If the system (1) uses the steering controller (89) with gain constraint (90), the scalar function (91) $V_{2}>0$ for all $[\tilde{\psi},\tilde{r}]^{\rm T}\neq\bm{0}$, the rate (92) $\dot{V}_{2}<0$ is negative definite except when $[\tilde{\psi},\tilde{r}]^{\rm T}=\bm{0}$, and the rate $\dot{V}_{2}\rightarrow 0$ as $t\rightarrow\infty$ locally around the orbit $O$. The value of the scalar function $V_{2}$ decreases along the trajectory of the system (1) to a lower bound. That lower bound is when $[\tilde{\psi},\tilde{r}]^{\rm T}=\bm{0}$—the system (1) is on the orbit $O$—because $\dot{V}_{2}\rightarrow 0$ as $t\rightarrow\infty$, and the rate $\dot{V}_{2}$ is only zero along the orbit $O$. Therefore, the system (1) is locally asymptotically attracted to the orbit $O$ when (90) is satisfied. ∎

### v2

@@ 5.1 Vector Field Guidance for the Desired Heading
where the “+” and “-” solutions are CCW and CW rotation, respectively. Expanding (58) with (63), (70), (71), and (72),
$\bm{u}$ $=\left(-\Phi^{\prime}+\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\right)\bm{e}_{r}\pm\left(\sqrt{v^{2}-\left(-\Phi^{\prime}+\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\right)^{2}}\right)\bm{e}_{\theta}\;.$ (73)
The vector field (73) consists of three parts shown in Fig. 6: (i) $-\Phi^{\prime}\bm{e}_{r}$ radial attraction, (ii) $(\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r})\bm{e}_{r}$ feed-forward, and (iii) circulation $\left(\sqrt{v^{2}-\left(-\Phi^{\prime}+\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\right)^{2}}\right)\bm{e}_{\theta}$.
Figure 6: The parts of a time-varying circulation vector field in (58) for counter-clockwise circulation: attraction, feedforward, circulation, and a combination of the attraction and feedforward, from left to right.
The combination of the radial attraction and the feed-forward terms $(-\Phi^{\prime}+\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r})\bm{e}_{r}$ are shown in Fig. 6 to compare the magnitude of the two radial fields. Trajectories of a particle moving in $\mathbb{R}^{2}$ using the controller $\bm{u}$ are visualized in Fig. 7. The vector field was created with a POI moving at $v_{g}=5$ m/s with heading $\gamma=-3\pi/4$ rad with $\beta=0.1$ and the UAV moving at $10$ m/s. The velocity streamlines do not arrive tangentially at the orbit because the orbit is translating to the bottom right corner—the UAV will not remain on the orbit in the next time step if the streamlines are tangent to the orbit.
Figure 7: The vector field (73) (gray), for an orbit (black) moving to the bottom left after $1$ s.
@@
$\bm{u}$ $=\left(\left(-v+v_{g}+\left|\dot{R}\right|\right)\frac{2}{\pi}\arctan(\beta(r-R))+\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R}\right)\bm{e}_{r}$
±$\pm\left(\sqrt{v^{2}-\left(\left(-v+v_{g}+\left|\dot{R}\right|\right)\frac{2}{\pi}\arctan\left(\beta(r-R)\right)+\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R}\right)^{2}}\right)\bm{e}_{\theta}\;.$ (74)
This controller builds upon the framework in [26] by finding a specific implementation of the generalized controller for a constant velocity single integrator model. To implement the controller, this work found functions $a_{1}$, $\Phi$, and $H$ that created a constant magnitude vector field that is tailored towards tracking the reference orbits introduced in Sec. 4.
@@ 5.2 Heading Rate
The vector field (74) written in terms of polar coordinates $r$ and $\theta$ is rotated to the rectangular coordinates with,
$\begin{bmatrix}u_{x}\\ u_{y}\end{bmatrix}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}u_{r}\\ u_{\theta}\end{bmatrix}\;.$ (75)
Since the magnitude of the vector field is the same as the Dubins vehicle speed, (31) rewritten in $xy$ coordinates is
@@
which is a function of the current $xy$ position and time $t$. If the the UAV starts with initial condition $\bm{q}_{0}=[q_{x},q_{y},\psi_{d}([q_{x},q_{y}]^{\rm T},0)]^{\rm T}$ then the system (1) with an unbounded turn rate can flow along the vector field $\bm{u}$ with the substitution $u_{\psi}=\dot{\psi}_{d}=v\kappa_{s}$ the expression of curvature from (3). Lemma 2 derives an expression for the curvature of a constant velocity curve in polar coordinates. By using the substitution $\dot{\bm{x}}_{\mathcal{I}}=\dot{\bm{\xi}}=\bm{u}$ with equation (20) from Lemma 2,
@@
where $\ddot{\bm{g}}=0$ since the POI has a constant velocity, and $\ddot{R}=0$ since the radius changes linearly. The derivative of (70) is
@@
$=\frac{(\dot{\xi}_{x}-\dot{g}_{x})(\xi_{x}-g_{x})+(\dot{\xi}_{y}-\dot{g}_{y})(\xi_{y}-g_{y})}{\sqrt{(\xi_{x}-g_{x})^{2}+(\xi_{y}-g_{y})}}\;,$ (81)
which can be simplified by substituting in (57), (11), and the polar direction $\bm{e}_{r}$ to give $\dot{r}=(\bm{u}-\dot{\bm{g}})\cdot\bm{e}_{r}=u_{r}-\dot{\bm{g}}\cdot\bm{e}_{r}$. Substituting (73) further simplifies this expression to
@@
$\dot{\Phi}^{\prime}$ $=\left(v-v_{g}-\left|\dot{R}\right|\right)\frac{2}{\pi}\frac{\beta(-[\left(v-v_{g}-\left|\dot{R}\right|\right)\frac{2}{\pi}\arctan\left(\beta(r-R)\right)])}{1+\beta^{2}(r-R)^{2}}$
@@
$\dot{\theta}$ $=\frac{(\bm{u}-\dot{\bm{g}})\cdot\bm{e}_{\theta}}{r}=\frac{u_{\theta}-\dot{\bm{g}}\cdot\bm{e}_{\theta}}{r}\;.$ (84)
@@
$\dot{\psi}_{d}$ $=\frac{\dot{\Phi}^{\prime}-\dot{\theta}\dot{\bm{g}}\cdot{{\bm{e}}_{\theta}}}{u_{\theta}}+\dot{\theta}=\frac{\dot{\Phi}^{\prime}}{u_{\theta}}+\dot{\theta}\frac{r}{u_{\theta}}\left(\frac{u_{\theta}-\dot{\bm{g}}\cdot\bm{e}_{\theta}}{r}\right)\;.$
@@
A plot of $\dot{\psi}_{d}$ is shown in Fig. 8
Figure 8: A surface plot and contour plots of $\dot{\psi}_{d}$ where the middle is cropped at $\dot{\psi}_{d}=4$ rad/s.
where the UAV’s speed is $10$ m/s, the POI moves with speed $5$ m/s in direction $\frac{-3\pi}{4}$, the gain $\beta=0.1$, and $t=0$ s. The value of $\dot{\psi}_{d}$ is small for all values outside of the orbit’s radius $10$ m, but there is a singularity where $\dot{\psi}_{d}$ goes to infinity at $r=0$ m. For large values of $\beta$ the sign is not strictly positive or negative.
@@ 5.3 Attraction to the Orbit
If the UAV is initialized with a heading $q_{\psi,0}\in\mathbb{S}$ that is not aligned with the vector field ${\bm{u}}$ then a controller is needed to drive it towards $\psi_{d}$. The following steering controller is proposed:
$u_{\psi}$ $=k_{\psi}\left(\psi_{d}-q_{\psi}\right)+\dot{\psi}_{d}|\left(\left(\bm{g}\cdot\bm{e}_{r}+\dot{R}\right)\frac{1-\cos(q_{\psi}-\psi_{d})}{q_{\psi}-\psi_{d}}+\frac{u_{\theta}\sin(q_{\psi}-\psi_{d})}{q_{\psi}-\psi_{d}}\right)\frac{\beta\frac{2}{\pi}\arctan\left(\beta(r-R)\right)}{1+\beta^{2}(r-R)^{2}}\;,$ (86)
where $\dot{\psi}_{d}$ is the rate of change of the heading for a trajectory flowing along the vector field $\bm{u}$ for $\bm{\xi}=[q_{x},q_{y}]^{\rm T}$, $k_{\psi}$ is a positive proportional gain, and the final term (derived in the next lemma) is required for Lyapunov-like stability [28].
@@
From [28, p. 125], a corollary of Barbalat’s lemma is the following: If a scalar function $V(\bm{q},t)$ satisfies, (i) $V(\bm{q},t)$ is lower bounded, (ii) $\dot{V}(\bm{q},t)$ is negative semi-definite, and (iii) $\dot{V}(\bm{q},t)$ is uniformly continuous in time, then $\dot{V}(\bm{q},t)\rightarrow 0$ as $t\rightarrow\infty$.
@@
where $\tilde{r}=r-R$ is the error between the UAV’s radial position and orbit $O$. The heading error $\tilde{\psi}=q_{\psi}-\psi_{d}$ is the error between the current and desired headings. The angular error $\tilde{\psi}\in\mathbb{S}$ is the smaller of the two distances from $\psi_{d}$ to $q_{\psi}$, traveling CW or CCW in $\mathbb{S}$. The function (88) has a lower bound of zero when $\tilde{\psi}=0$ and $\tilde{r}=0$. The derivative of (88) is
@@
$\dot{\tilde{r}}$ $=\dot{r}-\dot{R}=v\cos q_{\psi}\cos\theta+v\sin q_{\psi}\sin\theta-\dot{\bm{g}}\cdot\bm{e}_{r}-\dot{R}\;.$ (90)
@@
$\dot{\tilde{r}}$ $=v\cos\theta(\cos\psi_{d}\cos\tilde{\psi}-\sin\psi_{d}\sin\tilde{\psi})+v\sin\theta(\sin\psi_{d}\cos\tilde{\psi}+\sin\tilde{\psi}\cos\psi_{d})-\dot{\bm{g}}\cdot\bm{e}_{r}-\dot{R}\;.$ (91)
@@
$\dot{\tilde{r}}$ $=\cos\tilde{\psi}(u_{x}\cos\theta+u_{y}\sin\theta)-\sin\tilde{\psi}(-u_{x}\sin\theta+u_{y}\cos\theta)-\dot{\bm{g}}\cdot\bm{e}_{r}-\dot{R}\;.$ (92)
@@
$\dot{V}_{2}$ $=-k_{\psi}\tilde{\psi}^{2}-\frac{\left(v-v_{g}-\left|\dot{R}\right|\right)\cos(\tilde{\psi})\beta\frac{4}{\pi^{2}}\arctan^{2}{\beta\tilde{r}}}{1+\beta^{2}\tilde{r}^{2}}\;.$ (97)
@@
denote the coefficient appearing in (97). The bounds of $\epsilon$ were found using to be $0\leq\epsilon\leq\epsilon_{\rm max}$ where $\epsilon_{\rm max}$ is the solution to the equation $01-x\arctan{x}=$ 0, which is approximately $\epsilon_{\rm max}\approx 0.31483$. The function $\epsilon$ can be rewritten as
@@
$=\lim_{x\rightarrow\pm\infty}\frac{1}{2(1+x^{2})+4x^{2}}+\frac{1}{2(1+x^{2})}=0\;.$
@@
$k_{\psi}>v\beta\frac{4}{\pi^{4}}\epsilon_{\rm max}\geq\frac{\left(v-v_{g}-\left|\dot{R}\right|\right)\beta\frac{4}{\pi^{4}}\arctan^{2}{\beta\tilde{r}}}{1+\beta^{2}\tilde{r}^{2}}\;.$ (102)
Substituting $\tilde{\psi}=\pm\frac{\pi}{2}$ into (97), the inequality $\dot{V}<0$ leads to $k_{\psi}>0$. Local extrema occur where the derivative is zero:
$\frac{\partial}{\partial\tilde{\psi}}\dot{V}_{2}=0$ $=-2k_{\psi}\tilde{\psi}+\frac{\left(v-v_{g}-\left|\dot{R}\right|\right)\sin{\tilde{\psi}}\beta\frac{4}{\pi^{2}}\arctan^{2}{\beta\tilde{r}}}{1+\beta^{2}\tilde{r}^{2}}$
@@
If the system (1) uses the steering controller (86) with gain constraint (87), the scalar function (88) $V_{2}>0$ for all $[\tilde{\psi},\tilde{r}]^{\rm T}\neq\bm{0}$, the rate (89) $\dot{V}_{2}<0$ is negative definite except when $[\tilde{\psi},\tilde{r}]^{\rm T}=\bm{0}$, and the rate $\dot{V}_{2}\rightarrow 0$ as $t\rightarrow\infty$ locally around the orbit $O$. The value of the scalar function $V_{2}$ decreases along the trajectory of the system (1) to a lower bound. That lower bound is when $[\tilde{\psi},\tilde{r}]^{\rm T}=\bm{0}$ because $\dot{V}_{2}\rightarrow 0$ as $t\rightarrow\infty$, and the rate $\dot{V}_{2}$ is only zero along the orbit $O$. Therefore, the system (1) is locally asymptotically attracted to the orbit $O$ when (87) is satisfied. ∎